### v1

@@ Abstract
We present analyses of the early data from Rubin Observatory’s Data Preview 1 (DP1) for the globular cluster 47 Tuc field. The DP1 dataset for 47 Tuc includes four nights of observations from the Rubin Commissioning Camera (LSSTComCam), covering multiple bands (). We address challenges of crowding near the cluster core and toward the SMC in DP1, and demonstrate improved star-galaxy separation by fitting fifth-degree polynomials to the stellar loci in color-color diagrams and applying multi-dimensional sigma clipping. We compile a catalog of 3,576 probable 47 Tuc member stars selected via a combination of isochrone, Gaia proper-motion, and color-color space matched filtering. We explore the sources of photometric scatter in the 47 Tuc color-color sequence, evaluating contributions from various potential sources, including differential extinction within the cluster. Finally, we recover five known variable stars, including three RR Lyrae and two eclipsing binaries. Although the DP1 lightcurves have sparse temporal sampling, they appear to follow the patterns of densely-sampled literature lightcurves well. Despite some data limitations for crowded-field stellar analysis, DP1 demonstrates the promising scientific potential for future LSST data releases.
@@ 1 Introduction
As the earliest on-sky campaign, the Rubin Commissioning Camera (LSSTComCam) was used to align and verify the end-to-end functionality of the Rubin Observatory’s systems, resulting in 1,792 science-grade exposures collected in late 2024. These data were processed with the LSST Science Pipelines (Bosch et al., 2019; Developers, 2025) to produce Data Preview 1 (DP1)111https://dp1.lsst.io. The available DP1 data products comprise raw and calibrated single-epoch exposures, coadds, difference images, and various detection catalogs, totaling 3.5 TB. It contains 2.3 million distinct astronomical objects, and is accessible via the Rubin Science Platform, a cloud-based data science environment for data analysis. One of the DP1 target fields is centered on the globular cluster 47 Tucanae (47 Tuc).
@@
47 Tuc, aka NGC 104, is the second-brightest globular cluster in the sky, with mag. It’s distance was determined to be kpc using Gaia DR2 parallax by Chen et al. (2018), which corresponds to a true distance modulus of . The line-of-sight reddening toward this relatively high-latitude cluster was estimated at by Hesser & Philip (1976) based on Strömgren photometry of distant early-type stars along that sightline. This low value, consistent with its Galactic latitude of , has been propitious, making it possible for 47 Tuc to be used as a template for stellar populations in a metal rich globular cluster, whereas other metal rich ones are to be found in the Galactic plane and plagued with high reddening that varies on the scale of a cluster.
@@
Its Gaia distance is in agreement with the distance modulus of , derived from model analysis of a detached binary (V69) star in the cluster by Brogaard et al. (2017). Other prior values, such as from the modeling of color-magnitude diagram features, and the cooling curves of its white dwarfs are dependent on knowing or assuming the detailed metal content and the age of 47 Tuc’s stars. Thus Gaia’s parallax based distance sets the stage for a more sure-footed determination of the properties of the cluster’s stellar content, its age, and its history. Despite being an early known metal rich stellar system (e.g., (Rennó et al., 2020) or (Simunovic et al., 2023)), the morphology of its color-magnitude diagram implies an age of between and Gyrs (more specifically Thompson et al. (2020) claim Gyr), making it as old (to within discernible uncertainties) as the oldest globular clusters of the Galaxy. It has been studied using so called “chromosome mapping” of detailed metallicities, which show that its oldest stars likely followed multiple evolutionary and chemical enrichment pathways (Marino et al., 2023). Given its proximity, it also has been a target for studying the cooling sequence of its white dwarfs. Since 47 Tuc is a well-studied cluster, the parameter values listed here are illustrative examples rather than exhaustive or unique measurements.
@@
On the sky, 47 Tuc sits in front of the western extension of the Small Magellanic Cloud (SMC). A concentration of stars situated 15 from the center of 47 Tuc was first reported by Bellazzini et al. (2005) and surmised to be a background star cluster (Bologna A) and later confirmed to be an SMC cluster (Bellazzini et al., 2019). 47 Tuc itself has a highly concentrated core (Shapley class III), with claims that it contains an intermediate mass black hole (Della Croce, A. et al., 2024). By contrast, it also exhibits an extended periphery of stars: e.g. Piatti (2017) has claimed to find stars out to 5.5 times the tidal radius of , that indicate a halo like structure as seen in NGC 1851 (Olszewski et al., 2009). There is a radial gradient in the stellar population (Li et al., 2014), involving metallicities, helium abundance, and mass segregation. These characteristics collectively feed the speculation that like Cen and M54, 47 Tuc is the surviving central remnant from a dwarf galaxy merger.
@@
The complete 6-dimensional phase-space characterization of this globular cluster from Gaia provides details for good dynamical models of its motion through the Galaxy. If we are to use 47 Tuc as a laboratory for studying the detailed processes of its interactions within the Galactic field, it is necessary to trace the stars out as far as possible, but where stars belonging or originating in 47 Tuc must be tagged as distinct from the background of the SMC background stars; those from the ambient “stellar field” of the Galaxy through which 47 Tuc is passing; as well as the stars that may have been tidally stripped earlier from a possible dwarf galaxy progenitor of 47 Tuc. This study is an early look at real data from the Rubin Observatory systems which can inform us of the strengths and limitations of the data, as well as the state of our currently available methodology for attaining the required precision and accuracy in photometry and proper motions for disambiguating the stellar population components.
@@
This paper is organized as follows. In Section 2.2, we briefly describe the observation and the DP1 data products for the 47 Tuc field, and then present the field-specific data properties. Section 3 describes the 47 Tuc member star selection process, explores the photometry of selected member stars, and investigates the LSSTComCam detection of known variable stars. We summarize our conclusions and findings in Section 4.
@@ 2.2 DP1 Data Products for 47 Tuc
In full data release processing, all visit_image datasets are resampled onto a common sky projection (defined by a skymap), convolved to a common PSF, and then stacked (with inverse-variance weighting) into a “coadd” image. Outlier rejection is performed to prevent image artifacts, cosmic rays, and moving objects from printing through to the final coadd images. Object catalogs (NSF-DOE Vera C. Rubin Observatory, 2025b) are produced from detection and measurement of objects on the final deep_coadd images (NSF-DOE Vera C. Rubin Observatory, 2025c).
@@ 2.2.1 Depth and Coverage
Figure 2 shows the cumulative -band imaging depth (i.e., coadd depth), expressed as the 5 limiting magnitude, for the 47 Tuc field. This map effectively represents an idealized PSF magnitude limit, assuming perfect detection and measurement. The outlines of coadd patches are overlaid, with each tract distinguished by a unique color and line style. The bright core of 47 Tuc is saturated in the LSSTComCam commissioning data, leading to a lack of coadd coverage in that region. The absence of coadds in that region implies that no object measurements, difference imaging, difference-image source analysis, and forced photometry analysis could be performed.
To estimate the per-band areal coverage of the coadd images of the field, we identified valid regions in the depth map and traced their boundaries using a contour-based approach. This method captures the detailed shape of the coverage, including any internal gaps or holes. The resulting contours were converted to sky coordinates and used to construct a polygonal representation of the footprint. The total area was then computed from this geometry both including and excluding the saturated central region. The outermost boundary and the central hole determined for the band are shown in Figure 2. The outer boundary traces the full extent of the valid data, while the inner boundary delineates the saturated region excluded from the hole-corrected area measurement. The resulting sky coverage in the band is approximately 1.064 deg2 with the central hole included, and 0.915 deg2 when the hole is excluded. The areal coverage in other bands is listed in Table 1.
@@ 2.2.2 Impact of Crowding on Deblending
To understand the origin of the apparent lack of sources in the central and eastern parts of the field, we examine deblending outcomes for two representative tract/patch pairs: (454, 48), located on the eastern side where the resulting object density is low, and (453, 35), on the western side, which includes both dense and sparse regions. Figure 4 shows the spatial distribution of sources for which the deblending process was either performed or skipped due to excessive memory consumption or runtime requirements. In the eastern region (left panel), the notable drop in Object table source counts appears to be primarily driven by skipped deblending, likely due to crowding associated with the proximity to the SMC. Deblending can be skipped for various reasons, including excessively large parent footprints, too many detected peaks, a single-peak configuration, or excessive masking. In the western region (right panel), deblending was largely skipped in the crowded inner area but completed successfully in the outer parts, where the intrinsic object density is lower. This underscores the inherent challenges of performing photometry in crowded fields, a regime for which the LSST Science Pipelines are not specifically optimized. Improvements to scarlet (Melchior et al., 2018) lite should allow processing to push further into these and other dense regions before the first data release.
@@ 2.2.3 Star-Galaxy Separation
The LSST Science Pipelines report an extendedness parameter, which is based on the ratio of cModel and PSF magnitudes; point sources should have roughly equal fluxes from these two measures, while extended sources will have more flux in the model measurement than the PSF. Objects are classified as “not extended” (extendedness=0) if the ratio between their PSF and cModel fluxes satisfies . In this work, we use the refExtendedness parameter, which for each object corresponds to the extendedness measured in the “reference band” (selected as described in Sec. 3.4 of Bosch et al. 2018).
@@ 3.1.3 Selection in Color-color Diagrams
Figure 8 presents a summary of the sequential member selection process. The first three panels show color-magnitude diagrams illustrating the impact of each filtering step, from left to right: isochrone-based pre-selection, proper motion refinement, and final color-color diagram filtering. Each panel highlights the progressive narrowing of the stellar locus and removal of contaminants. The fourth panel compares the -band magnitude distribution of the initial isochrone-selected sample and the final sample after all filtering steps. The proper motion refinement effectively removed significant contamination from the SMC, particularly around its red clump, as well as brighter stars down to Gaia’s depth limit. In contrast, the color-color diagram filtering step primarily excluded faint objects—likely background galaxies—that could not be identified through Gaia proper motion due to their low brightness. The final sample consists of 3,576 member stars. The list of the selected 47 Tuc member stars is available online.
@@ 3.2.1 Saturation Limit
The brightest 47 Tuc member star (objectId: 579576081061783277) is on the red giant branch phase and has magnitudes of , , , and in the , , , and bands, respectively. The star has a Gaia counterpart, and its transformed photometry from the Gaia and to the LSSTComCam system (Tucker, 2025) agrees well within a few hundredth of a magnitude in the , , and , where no visit images have saturated pixels. In the band, however, the transformed magnitude differs by 0.06 mag, coinciding with mild saturation in the central pixels of a few visits. This offset is likely due to the fact that the band has the best seeing in the field (see Figure 1 and Table 1), leading to saturation in the core, and the RGB star peaks in flux near that band. The second brightest -band magnitude without saturated pixels is 15.19420.0002. In short, the unsaturated bright limits reported in the Object table for 47 Tuc are approximately 15.70, 15.08, 15.19, and 14.74 mag in the , , , and bands, respectively.

### v2

@@ Abstract
We present analyses of the early data from Rubin Observatory’s Data Preview 1 (DP1) for the globular cluster 47 Tuc field. The DP1 dataset for 47 Tuc includes four nights of observations from the Rubin Commissioning Camera (LSSTComCam), covering multiple bands (). We address challenges of crowding in the inner region of the cluster and toward the SMC in DP1, and demonstrate improved star-galaxy separation by fitting fifth-degree polynomials to the stellar loci in color-color diagrams and applying multi-dimensional sigma clipping. We compile a catalog of 3,576 probable 47 Tuc member stars selected via a combination of isochrone, Gaia proper-motion, and color-color space matched filtering. We explore the sources of photometric scatter in the 47 Tuc color-color sequence, evaluating contributions from various potential sources, including differential extinction within the cluster. Finally, of the 72 well-characterized variables in the field, we recover five known variable stars, including three RR Lyrae and two eclipsing binaries, in the coadd-based object catalog, and identify 62 in the difference image-based object catalog. Although the DP1 lightcurves have sparse temporal sampling, they appear to follow the patterns of densely-sampled literature lightcurves well. Despite some data limitations for crowded-field stellar analysis, DP1 demonstrates the promising scientific potential for future LSST data releases.
@@ 1 Introduction
As the earliest on-sky campaign, the Rubin Commissioning Camera (LSSTComCam) was used to align and verify the end-to-end functionality of the Rubin Observatory’s systems, resulting in 1,792 science-grade exposures collected in late 2024. These data were processed with the LSST Science Pipelines (Bosch et al., 2019; Developers, 2025) to produce Data Preview 1 (DP1111https://dp1.lsst.io; NSF-DOE Vera C. Rubin Observatory, 2025a). The available DP1 data products comprise raw and calibrated single-epoch exposures, coadds, difference images, and various detection catalogs, totaling 3.5 TB. It contains 2.3 million distinct astronomical objects, and is accessible via the Rubin Science Platform, a cloud-based data science environment for data analysis. One of the DP1 target fields is centered on the globular cluster 47 Tucanae (47 Tuc).
@@
47 Tuc, aka NGC 104, is the second-brightest globular cluster in the sky, with mag. Its distance has been measured with high precision to be 4.45 kpc using, for example, Hubble Space Telescope (HST) proper motion by Watkins et al. (2015) and Gaia DR2 parallax by Chen et al. (2018), which corresponds to a true distance modulus of . The line-of-sight reddening toward this relatively high-latitude cluster was estimated at by Hesser & Philip (1976) based on Strömgren photometry of distant early-type stars along that sightline. This low value, consistent with its Galactic latitude of , has been propitious, making it possible for 47 Tuc to be used as a template for stellar populations in a metal rich globular cluster, whereas other metal rich ones are to be found in the Galactic plane and plagued with high reddening that varies on the scale of a cluster.
@@
Its HST and Gaia distance is in agreement with the distance modulus of , derived from model analysis of a detached binary (V69) star in the cluster by Brogaard et al. (2017). Other prior values, such as from the modeling of color-magnitude diagram features, and the cooling curves of its white dwarfs are dependent on knowing or assuming the detailed metal content and the age of 47 Tuc’s stars. Thus Gaia’s parallax based distance sets the stage for a more sure-footed determination of the properties of the cluster’s stellar content, its age, and its history. Despite being an early known metal rich stellar system (e.g., (Rennó et al., 2020) or (Simunovic et al., 2023)), the morphology of its color-magnitude diagram implies an age of between and Gyrs (more specifically Thompson et al. (2020) claim Gyr), making it as old (to within discernible uncertainties) as the oldest globular clusters of the Galaxy. It has been studied using the so called “chromosome maps”, which trace abundance correlations among stellar populations and reveals that its oldest stars likely followed multiple evolutionary and chemical enrichment pathways (e.g., Carretta et al., 2005; Milone et al., 2012; Marino et al., 2023). Given its proximity, it also has been a target for studying the cooling sequence of its white dwarfs. Since 47 Tuc is a well-studied cluster, the parameter values listed here are illustrative examples rather than exhaustive or unique measurements.
@@
On the sky, 47 Tuc sits in front of the western extension of the Small Magellanic Cloud (SMC). A concentration of stars situated 15 from the center of 47 Tuc was first reported by Bellazzini et al. (2005) and surmised to be a background star cluster (Bologna A) and later confirmed to be an SMC cluster (Bellazzini et al., 2019). 47 Tuc itself has a highly concentrated core (Shapley class III), with claims that it contains an intermediate mass black hole (Della Croce, A. et al., 2024). By contrast, it also exhibits an extended periphery of stars: e.g. Piatti (2017) has claimed to find stars out to 5.5 times the tidal radius of , that indicate a halo like structure as seen in NGC 1851 (Olszewski et al., 2009). There is a radial gradient in the stellar population (e.g., Li et al., 2014), involving metallicities, helium abundance, and mass segregation. These characteristics collectively feed the speculation that like Cen and M54, 47 Tuc is the surviving central remnant from a dwarf galaxy merger.
@@
The addition of proper motions from Gaia for 47 Tuc completes the 6-dimensional phase space characterization of this cluster, providing the observational prerequisites for good dynamical models of its motion through the Galaxy (e.g., Katz et al., 2023). LSST will build on this foundation by delivering deeper, wider, and more time-resolved imaging, extending precise photometry and proper motions to far fainter magnitudes than Gaia can reach (Ivezić et al., 2019), improving constraints on the cluster’s internal and orbital kinematics. Such data are also essential for tracing 47 Tuc’s stellar populations to the largest possible radii and for cleanly separating its members from multiple sources of contamination: background stars from the SMC, ambient “stellar field” of the Galaxy through which 47 Tuc is passing, and any debris potentially stripped from a possible dwarf galaxy progenitor of 47 Tuc. DP1 provides an early look at real data from the Rubin Observatory systems which can inform us of the strengths and limitations of the data, as well as the state of our currently available methodology for attaining the required precision and accuracy in photometry and proper motions for disambiguating the stellar population components.
In this paper, we use the 47 Tuc field to evaluate DP1 performance in a particularly challenging regime and to demonstrate its early science potential. We aim to (1) showcase key DP1 data products for this field, (2) develop and apply methods for robust member-star selection and star-galaxy separation, (3) assess photometric precision, saturation limits, and sources of scatter, and (4) examine the recovery of known variable stars.
@@
This paper is organized as follows. In Section 2.2, we briefly describe the observation and the DP1 data products for the 47 Tuc field, and then present the field-specific data properties, such as crowding, deblending, and star-galaxy separation. Section 3 describes the 47 Tuc member star selection process, explores the photometry of selected member stars, and investigates the LSSTComCam detection of known variable stars. We summarize our main findings and briefly discuss potential with future data releases in Section 4.
@@ 2 Data for the 47 Tuc Field
In this section, we summarize the relevant DP1 observations for the 47 Tuc field, briefly describe the processing and data products, and highlight field-specific characteristics, such as depth, coverage, crowding, and star-galaxy separation performance.
@@ 2.1 Observations
We note that 47 Tuc lies within the footprint of the Wide-Fast-Deep (WFD) component of the LSST222https://survey-strategy.lsst.io/baseline/wfd.html. According to the current baseline strategy, each WFD field is expected to receive approximately 800 visits over the 10-year survey with the full LSSTCam (Lange et al., 2024; Roodman et al., 2024). While this DP1 dataset offers a valuable early look at Rubin performance in crowded fields, the final LSST dataset will deliver an order of magnitude more visits and significantly deeper coadds, improving both the photometric depth and temporal sampling on 47 Tuc.
@@ 2.2 DP1 Data Products for 47 Tuc
In full data release processing, all visit_image datasets are resampled onto a common sky projection (defined by a skymap333A skymap is a tiling of the celestial sphere that divides large-scale sky coverage into manageable sections for processing and analysis.), convolved to a common PSF, and then stacked (with inverse-variance weighting) into a “coadd” image. Outlier rejection is performed to prevent image artifacts, cosmic rays, and moving objects from printing through to the final coadd images. Object catalogs (NSF-DOE Vera C. Rubin Observatory, 2025b) are produced from detection and measurement of objects on the final deep_coadd images (NSF-DOE Vera C. Rubin Observatory, 2025c).
@@ 2.2.1 Depth and Coverage
Figure 2 shows the cumulative -band imaging depth (i.e., coadd depth), expressed as the 5 limiting magnitude, for the 47 Tuc field. This map effectively represents an idealized PSF magnitude limit, assuming perfect detection and measurement. The outlines of coadd patches are overlaid, with each tract distinguished by a unique color and line style. The DP1 skymap partitions the celestial sphere into 18,938 tracts, each spanning approximately 2.8 square degrees. Each tract is further subdivided into a 10 × 10 grid of uniformly sized patches, with each patch covering roughly 0.028 square degrees (NSF-DOE Vera C. Rubin Observatory, 2025a). The bright core of 47 Tuc is saturated in the LSSTComCam commissioning data, leading to a lack of coadd coverage in that region (see the hole in Fig. 2). The absence of coadds in that region implies that no object measurements, difference imaging, difference-image source analysis, and forced photometry analysis could be performed.
To estimate the per-band areal coverage of the coadd images of the field, we identified valid regions in the depth map and traced their boundaries using a contour-based approach. This method captures the detailed shape of the coverage, including any internal gaps or holes. The resulting contours were converted to sky coordinates and used to construct a polygonal representation of the footprint. The total area was then computed from this geometry both including and excluding the saturated central region. The outermost boundary and the central hole determined for the band are shown in Figure 2. The outer boundary traces the full extent of the valid data, while the inner boundary delineates the saturated region excluded from the hole-corrected area measurement. The resulting sky coverage in the band is approximately 1.064 deg2 with the central hole included, which spans 0.149 deg2, and 0.915 deg2 when the hole is excluded. The sizes of the central holes in the , , and bands are 0.155 deg2, 0.169 deg2, and 0.060 deg2, respectively. The areal coverage including and excluding the saturated region in other bands is listed in Table 1.
@@ 2.2.2 Impact of Crowding on Deblending
To understand the origin of the apparent lack of sources in the central and eastern parts of the field, we examine deblending outcomes for two representative tract/patch pairs: (454, 48), located on the eastern side where the resulting object density is low, and (453, 44), on the western side, which includes both dense and sparse regions.
Figure 4 shows the spatial distribution of sources for which the deblending process was either performed or skipped due to excessive memory consumption or runtime requirements. Deblending can be skipped for various reasons, including excessively large parent footprints, too many detected peaks, a single-peak configuration, or excessive masking.
In the eastern region (left panel, tract: 454, patch:48), the notable drop in Object table source counts appears to be primarily driven by skipped deblending, likely due to additional crowding associated with the proximity to the SMC. In such regions, the source detection algorithm correctly identifies peaks, but the majority of pixels in the patch are above the detection threshold, resulting in one giant detection footprint with a few small holes. This occurs when the footprints of many overlapping sources merge into a single, large blend, and the current implementation of scarlet lite skips these blends due to memory constraints. This failure mode arises not simply from stellar density, but from a combination of crowding and the brightness of overlapping sources, i.e., a few bright stars can trigger the same behavior as a dense cluster of faint ones. This underscores the inherent challenges of performing photometry in crowded fields, a regime for which the LSST Science Pipelines are not specifically optimized. Improvements to scarlet (Melchior et al., 2018) lite should allow processing to push further into these and other dense regions before the first data release.
In contrast, in the western region (right panel, tract: 453, patch:44), deblending was largely skipped in the crowded inner area but completed successfully in the outer parts, where the intrinsic object density is lower. The transition is pretty sharp, occurring almost vertically at RA = 4.85 deg. Unfortunately, measuring an accurate, representative crowding level at which the pipelines begin to struggle in DP1 is non-trivial due to complex spatial coverage variations and incompleteness that cannot be characterized without artificial star tests, which is beyond the scope of this paper.
@@ 2.2.3 Star-Galaxy Separation
The LSST Science Pipelines report an extendedness parameter, which is based on the ratio of composite model (cModel; Abazajian et al., 2004; Bosch et al., 2018) and PSF fluxes; point sources should have roughly equal fluxes from these two measures, while extended sources will have more flux in the model measurement than the PSF. Objects are classified as “not extended” (extendedness=0) if the ratio between their PSF and cModel fluxes satisfies . In this work, we use the refExtendedness parameter, which for each object corresponds to the extendedness measured in the “reference band” (selected as described in Sec. 3.4 of Bosch et al. 2018).
@@ 3 Analysis
In this section, we describe the membership determination procedure, including the improved star-galaxy separation described in Section 2.2.3, and examine the DP1 photometry for the resulting sample of cluster member stars. We also investigate the detection of variable stars in the 47 Tuc field.
@@ 3.1.3 Selection in Color-color Diagrams
Figure 8 presents a summary of the sequential member selection process. The first three panels show color-magnitude diagrams illustrating the impact of each filtering step, from left to right: isochrone-based pre-selection, proper motion refinement, and final color-color diagram filtering. Each panel highlights the progressive narrowing of the stellar locus and removal of contaminants. The fourth panel compares the -band magnitude distribution of the initial isochrone-selected sample and the final sample after all filtering steps. The proper motion refinement effectively removed significant contamination from the SMC, particularly around its red clump, as well as brighter stars down to Gaia’s depth limit. In contrast, the color-color diagram filtering step primarily excluded faint objects—likely background galaxies—that could not be identified through Gaia proper motion due to their low brightness. The final sample consists of 3,576 member stars. The list of the selected 47 Tuc member stars is available in a machine-readable format444https://zenodo.org/records/15794218. The catalog provides only the unique object IDs, that can be crossed matched to the DP1 data products via the Rubin Science Platform (RSP; Jurić et al., 2019; Economou, 2023).
@@ 3.2.1 Saturation Limit
The brightest 47 Tuc member star (objectId: 579576081061783277) is on the red giant branch phase and has magnitudes of , , , and in the , , , and bands, respectively. The star has a Gaia counterpart, and its transformed photometry from the Gaia and to the LSSTComCam system (Tucker, 2025) agrees well within a few hundredths of a magnitude in the , , and , where no visit images have saturated pixels. In the band, however, the transformed magnitude differs by 0.06 mag, coinciding with mild saturation in the central pixels of a few visits. This offset is likely due to the fact that the band has the best seeing in the field (see Figure 1 and Table 1), leading to saturation in the core, and the RGB star peaks in flux near that band. The second brightest -band magnitude without saturated pixels is 15.19420.0002. In short, the unsaturated bright limits reported in the Object table for 47 Tuc are approximately 15.70, 15.08, 15.19, and 14.74 mag in the , , , and bands, respectively.